OpenStudy (anonymous):

Jackson purchased a pack of game cards that was on sale for 22% off. The sales tax in his county is 6%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards.

OpenStudy (anonymous):

TheInspector

OpenStudy (anonymous):

put @ in front of my name to tag me

OpenStudy (anonymous):

im not sure on this one

OpenStudy (math&ing001):

Original price is y, so the price after taking 22% off it becomes : y1=y-y*22/100=y(1-22/100) Now after adding taxes, the price becomes y2=y1+y1*6/100 = y1(1+6/100) = y(1-22/100)(1+6/100) So y(1-22/100)(1+6/100) is the final price.

OpenStudy (anonymous):

The number of pages that Zev, Kelly, Marneisha, and Aleisha can read in a day is shown below. Zev read 15% of his 56-page book. Kelly read 12% of her 64-page book. Marneisha read 14% of his 68-page book Aleisha read 10% of her 72-page book. Who can read the greatest number of pages in a day?

OpenStudy (anonymous):

@TheInspector